OpenStudy (anonymous):

Solve for x. 7(x - 3) = 4(x + 5) x = 1/11 x = 1/3 x = 41/11 x = 41/3

OpenStudy (zzr0ck3r):

x = 41/3

OpenStudy (across):

Let\[7(x-3)=4(x+5).\]Then\[7x-21=4x+20,\]\[3x=41,\]\[x=\frac{41}{3}.\]

OpenStudy (waheguru):

7x-21=4x+20 7x-4x=20+21 x=413.

OpenStudy (waheguru):

\[x=41/3\]